You are given a chance to be a part of the Provincial Rescue Team for a day. While you

are on your first duty, the alarm goes off signaling that there is a fire outbreak at the

Robinsons Place Butuan. You and your team arrived at the place seeing a high-level fire

outbreak and noticed that there is an employee trapped on the rooftop which measures

43 m above the ground. You and your teammates were assigned to rescue the employee,

so you decided to let the employee jump from the rooftop to the trampoline you are

carrying. Your initial post is at the parking area, then you traveled alongside the building

1.20 m/s straight toward the employee. The trampoline you are carrying measures1.5 m

from the ground? If your team is to save the employee,

a. what must be the maximum time of your arrival directly to the bottom of the

employee?

b. where will you and your team position from your starting position?


1
Expert's answer
2021-04-28T13:03:21-0400

a) This will be the time it takes for the employee to reach the trampoline:


"t=\sqrt{\frac{2(H-h)}{g}}=2.9\text{ s}."

b) Moving at 1.2 m/s, you will cover a distance of


"d=vt=3.48\text{ m}."